### v1

# Evaluating Interventional Reasoning Capabilities of Large Language Models

Abstract
Numerous decision-making tasks require estimating causal effects under interventions on different parts of a system. As practitioners consider using large language models (LLMs) to automate decisions, studying their causal reasoning capabilities becomes crucial. A recent line of work evaluates LLMs ability to retrieve commonsense causal facts, but these evaluations do not sufficiently assess how LLMs reason about interventions. Motivated by the role that interventions play in causal inference, in this paper, we conduct empirical analyses to evaluate whether LLMs can accurately update their knowledge of a data-generating process in response to an intervention. We create benchmarks that span diverse causal graphs (e.g., confounding, mediation) and variable types, and enable a study of intervention-based reasoning. These benchmarks allow us to isolate the ability of LLMs to accurately predict changes resulting from their ability to memorize facts or find other shortcuts. Our analysis on four LLMs highlights that while GPT-4 models show promising accuracy at predicting the intervention effects, they remain sensitive to distracting factors in the prompts.
1 Introduction
Large language models (LLMs) based on Transformers demonstrate remarkable  [^1] Whether or not LLMs perform remarkably in an absolute sense is an active scholarly debate. Here, we mean that LLMs performance on tasks such as question-answering, summarization or serving as conversational agents is remarkable relative to earlier variants of statistical language models. performance on a variety of human-relevant tasks, from conversing to summarizing web-based information and answering complex question (Chen et al., 2021; Brown et al., 2020; Li et al., 2022; Katz et al., 2024). These breakthroughs have prompted an interest in integrating LLMs into decision-making. Consider two hypothetical scenarios:
Example 1 (Scientific discovery assistant)
A group of researchers provide an LLM with a text representation of a knowledge graph that is available in their discipline based on current scientific evidence. They want the LLM to generate candidate experiments based on putative cause-and-effect relationships. Importantly, as they come across new experimental conditions tried in the literature, they want the LLM to provide updated recommendations.
Example 2 (Automated A/B testing)
An e-commerce site runs A/B tests on features of their site, collecting relevant engagement metrics as outcomes. They want an LLM to process these tabular datasets that record which experimental condition each user was assigned to and what their outcome was, and determine which feature value between A and B to adopt. Importantly, they plan to re-query the LLM habitually based on evidence from newly run A/B tests.
Both scenarios require reasoning about how new experimental conditions generated by interventions change the state of our knowledge. Example 1 involves assessing which causal links become plausible upon receiving information about interventions that were performed. Example 2 requires knowing how interventions – new A/B tests in this case – change the causal conclusions we previously drew. In these examples, to support decision-making, LLMs need to integrate information about interventions and appropriately update their beliefs about the the data-generating model they have.
Contributions. Motivated by the role that the calculus of interventions plays in decision-making, this paper makes the following contributions:
  1. 1.
    We introduce intervention effect (IE) prediction, where, given a causal graph and an intervention on a variable in that graph, one must determine how the directed path between two specific variables changes. In essence, IE prediction provides a concrete task that tests the ability of LLMs to appropriately update their beliefs after receiving information about an intervention experiment.
  2. 2.
    To assess the degree to which LLMs can accurately reason about interventions, we propose a methodology to turn instances of these classification problems into a prompt and introduce an intervention effect prediction benchmark.
  3. 3.
    Recognizing the sensitivity that LLMs show to prompt-design, we design studies to further disentangle the effects on IE prediction performance due to spurious properties of the prompt like the names of variables and the causal relations they form.
This paper contributes to the emerging body of work on evaluating algorithmic causal reasoning in LLMs (Cai et al., 2023; Jin et al., 2023; 2024; Binz & Schulz, 2023). In Section 5, we summarize the connections to several threads of related work.
The empirical studies we conduct on four LLMs suggest that, under certain choices for generating prompts based on intervention effects, variants of GPT-4 achieve promising accuracy of predicting what happens to causal graph relationships under interventions. However, we find that LLMs are sensitive to the choice of variable names when we instantiate prompts. When prompts contain facts that LLMs might have been trained on, their performance drops significantly. This result underscores the importance of careful design when creating benchmarks to avoid drawing spurious conclusions.
2 Preliminaries
We summarize the aspects of large language models (LLMs) and causality that are necessary to understand the empirical study we design in this paper.
Large language models.  For our purposes, it suffices to view LLMs (Touvron et al., 2023; OpenAI, 2023) as models of conditional distributions $P_{\theta}(Y=y|x)$ where $x$ is a sequence of tokens called context, the random variable $Y$ is the next item in the sequence and $y$ is a token (the smallest unit that text was broken down to for LLM processing). The parameter vector $\theta$ captures all relevant LLM parameters such as the self-attention and multi-layer perceptron (MLP) weights at each layer. In this paper, we focus on inference, the task of evaluating $P_{\theta^{*}}(Y=y|x)$ on unseen contexts $x$ given an LLM with trained parameters $\theta^{*}$  [^2] The parameters $\theta^{*}$ approximately maximizes the log probability $P_{\theta}(Y=y|x)$ on web-scale text.. Specifically, we design benchmarks where the contexts pose yes/no causal reasoning questions and we evaluate the output produced by the LLM, generated by $y\sim P_{\theta^{*}}(Y=y|x)$.
Causal directed acyclic graphical models and perfect interventions.  Causal directed acyclic graphical models (causal DAGs) provide a way to encode a particular class of interventions using minimal changes to the graph structure (Peters et al., 2017). More technically, in a causal DAG $G=(V,E)$ over variables $V$ with directed edges in the set $E$, we represent a perfect intervention $\mathrm{do}(V_{i}=v)$ that sets the value of variable $V_{i}$ to $v$ by modifying the DAG $G$ to delete the edges going into $V_{i}$.
While causal DAGs and perfect interventions do not allow us to express all possible causal questions, we focus on them in this paper because presence/absence of edges are well-suited to constructing binary classification task, which simplify LLM evaluation by enabling us to evaluating the probability of just two tokens, yes and no, rather than needing to evaluate arbitrary utterances.
Figure 1: Causal DAGs. In the empirical studies, we define intervention effects based on three causal DAGs: bivariate, confounding and mediation graphs.
Understanding how interventions modify the data-generating process is key to knowing which causal inferences a dataset can (or cannot) support (Pearl, 2009; Peters et al., 2017). To understand this distinction, consider the confounding DAG in the middle in Figure 1. Suppose this graph represents the beliefs of a scientist who is trying to reason about a possible causal relationship between variables $B$ and $C$. In the absence of experimental data, they know that their inference is hopeless since a third variable $A$, called a confounding variable or confounder, can equally well explain all of the association observed between the variables $B$ and $C$. However, if they learned that an experiment had been conducted to intervene on $B$, the scientist knows that associations between $B$ and $C$ would now capture the causal effect of $B$ on $C$, since the intervention removed the variable $A$’s confounding influence on $B$. If, unlike human experts, LLMs cannot reliably predict the graphical changes made by intervening, they run the risk of drawing invalid conclusions for causal questions that are in fact non-identifiable given available evidence.
In Section 4, we design empirical studies based on the bivariate DAG in the left of Figure 1, the confounding DAG in the center in Fig. 1, and the mediation DAG in the right in Fig. 1, where a third variable $B$ mediates the effect that $A$ has on $C$. In the bivariate DAG, intervening on $B$ severs the link between $A$ and $B$ and in the mediation DAG, intervening on $B$ similarly means that the variable $A$ can no longer affect $C$.
3 Methodology: Defining the causal reasoning task
In this section, we introduce and define intervention effects, which quantify the changes to graphical relations entailed by a causal DAG $G$ under different $\mathrm{do}$-interventions.
3.1 Intervention effect
To formalize effects that interventions have on causal DAGs, we need to define graphical features of DAGs that are of interest.
Definition 1 (Causal relation)
Given a causal DAG $G=(V,E)$, a causal relation $C_{uv}(G)$ involving variables $V_{u}$ and $V_{v}$ in DAG $G$ is 1 if there is a directed path $e_{ua}\rightarrow\ldots\rightarrow e_{bv}$ in $G$ from $V_{u}$ to $V_{v}$ (where each edge $e_{ij}$ along the path is in $G$) and 0 otherwise.
In words, a causal relation captures whether or not a variable exerts an indirect or direct causal influence on another variable in a particular DAG $G$ that contains these variables. In this paper, we focus on causal relations as a way to assess how interventions affect causal DAG structure.
With causal relations in hand, we now define the main quantity of interest in this paper, the intervention effect (IE).
Definition 2 (Intervention effect)
Given a causal DAG $G=(V,E)$, a variable $V_{i}$ on which we perform a perfect intervention captured by $\mathrm{do}(V_{i}=*)$ (where we use “*” to indicate that we do not care about the value that $V_{i}$ is set to), and a query causal relation $C_{uv}$, an intervention effect (IE) is,
$\kappa^{G}_{i}(C_{uv})=C_{uv}(G)-C_{uv}(G^{i}),$ (1)
where the DAG $G^{i}$ is the modification of DAG $G$ under an intervention to the variable $V_{i}$.
An IE $\kappa^{G}_{i}$, which is defined relative to a base DAG $G$ and intervention target $V_{i}$, is a function of a particular causal relation $C_{uv}$ of interest. Since interventions only require us to remove the edges incoming to $V_{i}$, note that we expect many intervention effects $\kappa^{G}_{i}(C_{uv})$ to evaluate to 0, indicating that the intervention left at least one directed path between the variables $V_{u}$ and $V_{v}$ unaffected.
3.2 LLM predictor of intervention effects
In this paper, we ask the question: how can we leverage LLMs to predict intervention effects (IEs), given that they take textual contexts as input? To evaluate LLMs on this prediction task, we need to instantiate mapping functions,
$\begin{split}m_{G}:G\rightarrow x_{G};\quad m_{I}:V_{i}\rightarrow x_{i};\quad m_{C}:C_{uv}(G)\rightarrow x^{G}_{uv},\end{split}$ (2)
which take a DAG $G$, an intervention target $V_{i}$ and a causal relation $C_{uv}(G)$ in DAG $G$ as input and output text sequences $x_{G}$, $x_{i}$ and $x^{G}_{uv}$ to verbalize these facts as inputs to an LLM. Additionally, the concatenation function $c(x_{G},x_{i},x^{G}_{uv})$ produces a coherent context $x$ while keeping the information in the input sequences intact.
In the empirical studies in Section 4, we instantiate these mapping functions concretely, but for now, assume that these mapping functions are given. To ease notation later, we will define one more function,
$O(x)=\mathbb{1}[y\sim P_{\theta}(Y=y|x)=\texttt{yes}],$ (3)
which returns 1 when the next word sampled by the LLM given the context $x$ is yes.
An LLM-based predictor of the intervention effect is a function,
$\hat{\kappa}^{G}_{i}(C_{uv})=O(c(x_{G},x^{G}_{uv}))-O(c(x_{G},x_{i},x^{G_{i}}_{uv})).$ (4)
In words, we turn LLM into predictors by contrasting their responses to two different sequences $x$ and $x_{0}$. In the sequence $x_{0}=c(x_{G},x^{G}_{uv})$, we describe a causal DAG $G$, describe no intervention, and ask about a causal relation in this base graph $G$. In the other sequence $x=c(x_{G},x_{i},x^{G_{i}}_{uv})$, we describe the same causal DAG $G$, but now, introduce the description of an intervention on some target $V_{i}$, and ask about the same causal relation but in the modified graph $G^{i}$. Note that we never describe the complete graph post intervention ($x_{G_{i}}$) in the sequence $x$. Hence, the LLM has to infer the $G_{i}$ and subsequently use it to reason about the causal relation post intervention ($C_{uv}(G_{i})$).
3.3 Accuracy metric for LLM prediction
With a concrete question – intervention effects – in mind and a strategy for producing predictions of intervention effects using LLMs, what remains is defining a metric that measures the fidelity of LLM predictions. That is, we want to compare an estimated IE $\hat{\kappa}^{G}_{i}(C_{uv})$ against the ground-truth IE $\kappa^{G}_{i}(C_{uv})$.
The straightforward and natural choice for such a comparison would be accuracy,
$\mathrm{acc}(G,i,u,v)=\mathbb{1}[\kappa^{G}_{i}(C_{uv})=\hat{\kappa}^{G}_{i}(C_{uv})].$ (5)
However, this accuracy metric can return 1 scenarios such as,
$C_{uv}(G)=1,\,\,C_{uv}(G^{i})=1,\,\,\hat{C}_{uv}(G)=0,\,\,\hat{C}_{uv}(G^{i})=0,$ (6)
where (with some slight abuse of notation) $\hat{C}_{uv}(\cdot)$ denotes the LLM’s prediction about whether or not a causal relation is true in a given DAG. In this example, a target causal relation $C_{uv}$ is true in both the base causal DAG $G$ and its modified, post-intervention counterpart $G^{i}$, but the LLM incorrectly predicts that these causal relations are false under both graphical scenarios. The accuracy metric misleads us when the LLM correctly predicts that the causal relation does not vary, but does not parse the causal relation correctly in either graphical scenario. Thus, to ensure that the accuracy is 0 in such cases, we slightly modify the accuracy metric to be,
$\mathrm{acc}(G,i,u,v)=\mathbb{1}[\kappa^{G}_{i}(C_{uv})=\hat{\kappa}^{G}_{i}(C_{uv})]\cdot\mathbb{1}[C_{uv}(G)=\hat{C}_{uv}(G)].$ (7)
4 Experiments
We introduce benchmarks that verbalize intervention effects as prompts to LLMs. Using these benchmarks, we empirically investigate four research questions: 1) how accurately do LLMs predict causal relation changes under interventions to causal DAGs on benchmarks generated with randomly chosen variable names? 2) to what extent is LLM performance affected when the instantiated causal relations reflect potentially memorized facts? 3) could LLMs be predicting the effects of interventions accurately by learning a shortcut instead? 4) How robust are LLMs to the way interventions are described in-context?
We find that while some LLMs demonstrate good performance on the first benchmark, if the prompts are designed so that some textual graph relationships are well-known causal facts, the performance can drop, pointing to the importance of stress-testing LLMs when evaluating them for abilities such as causal reasoning. We provide the full experimental details in Appendix A and additional experiment results  [^3] Code for experiments: https://github.com/tejaskasetty/interventional-reasoning-llms in Appendix B.
Figure 2: An illustration of the mapping functions that verbalize the information for a single intervention effect estimation task into a prompt.
4.1 Dataset Generation
In this paper, we design three benchmarks, each consisting of three scenarios that correspond to the three causal DAGs in Figure 1. In each scenario, we generate intervention effects by considering an intervention on each variable in the DAG, and enumerating over all causal relations in the DAG per intervention target. This corresponds to 22 intervention effects across all scenarios.
To create prompts for each intervention effect, we need to define the mapping functions in Equation 2. Figure 2 illustrates the template we use to map DAGs, interventions, as well as causal relations to generate prompts. To instantiate these templates, we need to sample names for variables in the DAG.
The three benchmarks differ in how these names are sampled, with each benchmark designed to help us investigate a different research question. We summarize the differences across benchmarks:
  1. 1.
    Random ($\mathcal{D}_{1}$): Variable names are randomly chosen English characters.
  2. 2.
    Tübingen ($\mathcal{D}_{2}.$): Variables names are chosen from entities that appear in the Tübingen pairs (TP) dataset of causal relations (Mooij et al., 2016).
  3. 3.
    Anti-commonsense ($\mathcal{D}_{3}$): Variable names are chosen so that some of the causal relations are the opposite of those found in TP.
For each benchmark and intervention effect therein that we want to turn into a prompt, we sample variable names fifteen times to have multiple independently generated examples per intervention effect (allowing us to report significant differences later).
4.2 Empirical studies on benchmarks
We investigate the four research questions (RQs) on the proposed benchmarks, studying four LLMs: GPT-3.5, GPT-4, GPT-4-turbo (OpenAI, 2023), and LLaMA-2 (Touvron et al., 2023). For reporting the IE prediction accuracy, unless mentioned otherwise, we aggregate the performance over different causal relation queries using mean for each scenario (causal graph, intervened variable). Further, we always report the mean $\pm$ standard error across 15 choices for the prompt entities.
RQ1: How accurate are LLMs at predicting the effects of interventions?
Graph Type Bivariate Confounding Mediation
Intervened Variable A B A B C A B C
GPT-3.5 $0.83\pm 0.08$ $0.87\pm 0.06$ $0.80\pm 0.09$ $0.69\pm 0.12$ $0.36\pm 0.09$ $0.58\pm 0.11$ $0.36\pm 0.12$ $0.67\pm 0.12$
GPT-4 $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.78\pm 0.09$ $0.82\pm 0.08$ $0.96\pm 0.03$
GPT-4-turbo $1.0\pm 0.0$ $0.97\pm 0.03$ $0.96\pm 0.03$ $0.93\pm 0.05$ $1.0\pm 0.0$ $0.98\pm 0.02$ $1.0\pm 0.0$ $1.0\pm 0.0$
LLaMA-2 $0.50\pm 0.12$ $0.40\pm 0.12$ $0.56\pm 0.12$ $0.53\pm 0.11$ $0.16\pm 0.06$ $0.69\pm 0.09$ $0.56\pm 0.12$ $0.64\pm 0.12$
Table 1: IE prediction accuracy on the Random $(\mathcal{D}_{1})$ benchmark. GPT-4 variants are the best performing models, while LLaMA-2 appears to struggle with interventional reasoning.
To study this first RQ, we evaluate the performance of the four LLMs on $\mathcal{D}_{1}$, the Random benchmark, measuring their accuracy at predicting intervention effects using the metric we introduced in Equation 7. Table 1 summarize these results. We see that GPT-based models perform notably better than LLaMA-2, with GPT-4-turbo demonstrating near-perfect accuracy across all effects. LLaMA’s performance suggests that it is not a reliable model for predicting the impacts of interventions, at least in a zero-shot way. In what follows in the main paper, we focus on results for GPT models, deferring LLaMA results to the appendix (Section B).
Model Graph Intervened Variable Tübingen ($\bf\mathcal{D}_{2}$) Anti-commonsense ($\bf\mathcal{D}_{3}$)
GPT-3.5 Bivariate B $0.83\pm 0.10$ $0.63\pm 0.12$
Mediation B $0.33\pm 0.12$ $0.40\pm 0.13$
C $0.67\pm 0.12$ $0.67\pm 0.12$
GPT-4 Bivariate B $1.0\pm 0.0$ $0.97\pm 0.05$
Mediation B $1.0\pm 0.0$ $0.73\pm 0.11$
C $1.0\pm 0.0$ $1.0\pm 0.0$
GPT-4-turbo Bivariate B $0.90\pm 0.08$ $0.57\pm 0.13$
Mediation B $1.0\pm 0.0$ $1.0\pm 0.0$
C $1.0\pm 0.0$ $1.0\pm 0.0$
Table 2: IE prediction performance on specific scenarios for benchmarks ($\mathcal{D}_{2},\mathcal{D}_{3}$) to understand the role memorization. Even best performing models in Table 1 like GPT-4 & GPT-4-turbo are negatively impacted by memorization effects in the Mediation scenario.
RQ2: To what extent is LLM performance affected by possibly memorized causal relations?  While we might be tempted to conclude that GPT-4 reliably predicts changes to models after interventions are performed, we consider spurious factors that can affect model performance. In particular, Kıcıman et al. (2023) found that GPT models reliably retrieved information about TP causal relations, suggesting that these relations could have been included in the training data for LLMs. (We reproduce their findings in Appendix Table 6.) This leads to a worrying possibility: after interventions, could LLMs fail to update their beliefs about causal relations that they have potentially memorized?
To study this RQ, we consider the anti-commonsense benchmark, $\mathcal{D}_{3}$. We consider the intervention effects defined based on: 1) for the bivariate DAG in Figure 1, consider an intervention on $B$ and consider the causal relations $C_{AB}$; for the mediation DAG in Figure 1, consider interventions on $B$ and on $C$, and causal relations $C_{AC}$. Crucially, these intervention effects were chosen so that if an LLM attends only to the plausibly memorized causal relation, e.g., $C_{AB}(G)$ in the bivariate DAG $G$, then under intervention to $B$, where the causal relation $C_{AB}(G^{B})$ changes, the LLM will be penalized if it relies on its memorized fact. This reasoning holds true for intervention effects defined the mediation DAG as well.
We contrast the performance that LLMs achieve on this anti-commonsense benchmark against their performance on the Tübingen benchmark, $\mathcal{D}_{2}$. In $\mathcal{D}_{2}$, although the names of variables come from the same distribution (the TP dataset), the target causal relations such as $C_{AB}$ are not ground-truth relations in TP, meaning that LLMs cannot rely on their memory when being evaluated on this benchmark. If we saw no statistically significant differences in prediction when contrasting the same exact intervention effect but with prompts instantiated according $\mathcal{D}_{1}$ versus $\mathcal{D}_{2}$, then we would have stronger evidence that LLMs might not be distracted by the presence of plausibly memorized causal relations. Table 2 summarizes the results across the two benchmarks. We see that the previously top-performing GPT models are all affected by distracting causal relations when considering effects in the bivariate DAG, with GPT-4-Turbo’s accuracy decreasing to 55%. Interestingly, these models fare better in the mediation DAG setting. However, these findings point to the significant impacts that distracting aspects of prompts can have on LLM reasoning, suggesting continued caution if using LLMs for decision-making tasks.
RQ3: Could LLMs be learning a shortcut to predict intervention effects?  Consider the confounding DAG (Figure 1) and the causal relation $C_{BC}(G)$ which does not change under intervention on the variable $A$. In these examples, LLMs that accurately parse causal relations from text descriptions of DAGs would also obtain accurate IE estimates. Hence, predicting causal relations from the input graph in text – call this task relation retrieval – offers a shortcut: an LLM can attend to tokens in the context to solve relation retrieval and still perform good at IE estimation, thereby confounding the conclusions that can be drawn with this benchmark.
However, the intervention effects we defined to study RQ2 offer an insight into how we can disentangle relation retrieval and accurate IE prediction. Notice that the IEs we defined to study RQ2 characterize scenarios where causal relations differ between the base DAG $G$ and the post-intervention DAG. Thus, LLMs that rely only on relation retrieval cannot accurately estimate IEs. Building on this insight, we divide all the IEs $\kappa^{i}_{G}(C_{uv})$ into two groups:
  1. 1.
    $\bf IE=0$: Graph doesn’t change as a result of intervention; $C_{uv}(G)=C_{uv}(G^{i})$.
  2. 2.
    $\bf IE=1$: Graph changes as a result of intervention; $C_{uv}(G)\neq C_{uv}(G^{i})$
Graph Type Bivariate Confounding Mediation
Intervened Variable B B C B C
GPT-3.5 IE = 0 $0.93\pm 0.06$ $0.6\pm 0.13$ $0.47\pm 0.13$ $0.33\pm 0.12$ $0.67\pm 0.12$
IE = 1 $0.8\pm 0.1$ $0.67\pm 0.12$ $0.2\pm 0.1$ $0.37\pm 0.12$ $0.67\pm 0.12$
GPT-4 IE = 0 $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.87\pm 0.09$ $1.0\pm 0.0$
IE = 1 $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.8\pm 0.1$ $0.93\pm 0.06$
GPT-4-turbo IE= 0 $1.0\pm 0.0$ $0.96\pm 0.05$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
IE = 1 $0.93\pm 0.06$ $0.93\pm 0.06$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
Table 3: IE prediction performance across sub-cases to isolate the effect of relation retrieval on Random $(\mathcal{D}_{1})$ benchmark. LLMs do not significantly rely on shortcuts related to relation retrieval from the input prompt. Since intervening on variable A never changes the causal graph in any scenario, we do consider them for this analysis and leave them blank.
Note that drop in performance on group ($IE=1$) as compared to the group ($IE=0$) can indicate reliance on shortcuts based on relation retrieval. We consider the average performance of the LLMs on the effects in each group, selecting only those LLMs which achieve an accuracy $\geq 0.95$ on relation retrieval (which we report in Appendix Table 9). We focus on the random benchmark ($\mathcal{D}_{1}$) to exclude any impacts due to memorized variable names. Table 3 summarizes this study. We find that the general trend does not show strong reliance on shortcuts across LLMs; only GPT-3.5 has a significant drop in its relative performance on group ($IE=1$) vs group ($IE=0$) for the confounding DAG case.
Graph Type Bivariate Confounding Mediation
Intervened Variable A B A B C A B C
GPT-3.5 $0.67\pm 0.12$ $0.83\pm 0.06$ $0.56\pm 0.11$ $0.58\pm 0.11$ $1.0\pm 0.0$ $0.62\pm 0.10$ $0.44\pm 0.10$ $0.4\pm 0.12$
GPT-4 $0.97\pm 0.03$ $1.0\pm 0.0$ $0.89\pm 0.05$ $0.82\pm 0.07$ $1.0\pm 0.0$ $0.96\pm 0.03$ $0.76\pm 0.10$ $1.0\pm 0.0$
GPT-4-turbo $0.9\pm 0.07$ $0.97\pm 0.03$ $0.96\pm 0.03$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
Table 4: IE prediction accuracy on the Random $(\mathcal{D}_{1})$ benchmark for the substitution task. The performance of GPT-3.5 & GPT-4 is worse under substitution compared to the non-substitution case (Table 1), while GPT-4-turbo does not show significant change.
RQ4: Are LLMs robust to descriptions of interventions in-context?  Consider the mapping function $M_{I}(\cdot)$ that verbalizes an intervention in Figure 2. We ask whether LLMs can achieve the same performance on the intervention effects in the random benchmark if we varied the mapping function $M_{I}(\cdot)$ to instead “teach” an LLM (Patel et al., 2023) about a new graphical operation that behaves identically to an intervention. We randomly generate strings to instantiate this operation. Figure 3 in the appendix illustrates how we generate prompts for this task which we refer to as the substitution task. Table 4 summarizes IE prediction accuracy for the substitution task on the Random benchmark $\mathcal{D}_{1}$. Contrasting these results against Table 1, we see that for GPT-4 and GPT-3.5, the performance generally suffers. However, GPT-4-turbo appears to be robust to changes in the way interventions are described.
5 Related Work
This paper relates most closely to recent papers that develop benchmarks to evaluate LLMs on various causal reasoning tasks. Kıcıman et al. (2023) introduced multiple causal reasoning benchmarks for LLMs, including evaluating the ability of LLMs to recover the bivariate causal DAGs introduced in the Tübingen pairs dataset (Mooij et al., 2016). Kıcıman et al. (2023) found that GPT models recovered known causal relationships with up to 96% accuracy when experimenting with various prompting strategies such as including system prompts. Building on these results, several recent papers use LLMs to augment the discovery of causal DAGs using natural language descriptions of variables (Abdulaal et al., 2024; Castelnovo et al., 2024; Jiralerspong et al., 2024; Takayama et al., 2024; Long et al., 2023b; a; Ban et al., 2023; Vashishtha et al., 2023). However, evaluating LLMs on their ability to retrieve causal knowledge about known variables constitutes commonsense causal reasoning. In contrast, this paper contributes to work that evaluates abstract causal reasoning (Binz & Schulz, 2023; Jin et al., 2024; 2023; Cai et al., 2023; Liu et al., 2024), assessing the ability of LLMs to use axioms of causality to solve tasks involving general or even new variables.
In the vein of causal reasoning, Jin et al. (2023) studied whether LLMs could correctly infer some causal relationships based on conditional independence statements, comparing LLM predictions to those made by an oracle causal discovery algorithm for observational data, where all causal relations cannot be resolved. In contrast to observational causal discovery, this paper focuses on reasoning about interventions. Also focusing on interventions, Jin et al. (2024) introduced CLadder, a comprehensive benchmark that includes the estimation of causal effects from quantitative data. Causal effect estimation is a complex task that requires solving multiple sub-tasks such as: (i) parsing the prompt to extract a causal DAG, (ii) inferring a function that estimates the effect given the DAG, and (iii) applying that function to the given quantitative data. Concurrently, Cai et al. (2023) introduced a task that asks LLMs to output only causal relationships given a tabular dataset that includes variable names. They focus on disentangling the impacts that prior knowledge (e.g., variables names) and quantitative data have on LLM performance. The empirical study we conduct to assess whether LLMs are sensitive to the presence of plausibly memorized causal relations is similar to experiments conducted by Cai et al. (2023). In contrast to these benchmarks, intervention effects target a narrower question than the general estimation of causal effects, since intervention effects involve binary classification only (i.e., the absence/presence of causal relations in DAGs). We argue that the evaluations we design better isolate causal reasoning from sub-tasks like drawing statistical inferences from quantitative data provided in-context, which both CLadder and the work of Cai et al. (2023) require.
In focusing on intervention effects, we build on the work of Binz & Schulz (2023), who were motivated by prior work in psychology (Waldmann & Hagmayer, 2005) that shows humans weight collected observational evidence and experimental evidence differently when drawing causal conclusions. Binz & Schulz (2023) adapted this psychology study for LLMs, creating prompts that describe observational and post-interventional findings to LLMs to see if they update their beliefs about a system after interventions. They found that GPT-3, unlike human subjects, fared poorly at understanding the implications of interventions. Motivated by their focus on intervention-based reasoning, we significantly expand on the evaluations designed by Binz & Schulz (2023), systematically generating intervention effects with varying degrees of difficulty to further explore the effects of plausible memorization and shortcuts like relation retrieval.
Finally, we note that recent work (Zečević et al., 2023) has called into question whether LLMs are causal parrots, mimicking behavior seen during training without being able to generalize the use of causal logic. Kıcıman et al. (2023); Cai et al. (2023); Jin et al. (2024) as well as the benchmarks designed in this paper aim to stress-test LLMs more extensively by varying prompt generation to assess different shortcuts that LLMs could have learned.
6 Discussion and Limitations
The goal of this paper is to introduce a causal reasoning benchmark that stress-tests the ability of LLMs to accurately predict how knowledge should be updated after interventions are performed, without conflating other aspects of reasoning such as statistical inference on quantitative data. The research questions that we investigated point to both some optimism and caution. While on one hand, in some scenarios, GPT-4 appears to accurately predict how interventions modify given causal relations, on the other hand, its performance can be negatively affected when prompts describe causal knowledge that it has plausibly memorized from training. Overall, these findings point to the continued importance of designing benchmarks and studies that evaluate varied aspects of abstract causal reasoning in LLMs, especially if practitioners wish to use LLMs to generate candidate decisions.
While the intervention effect prediction task we define in this paper has the benefit of being easy to evaluate, since it requires binary responses, the findings that this task can suggest are also limited. For example, IE prediction cannot help us assess how accurately LLMs perform causal identification, the process of deciding which causal inferences can be made given a causal DAG. Moreover, we focus on evaluation in this paper and do not propose methods for improving causal reasoning in LLMs via few-shot learning or fine-tuning. Both of these limitations point to future research directions that we think are worth exploring.
Finally, keeping in mind Goodhart’s law 444en.wikipedia.org/wiki/Goodhart’s_law, we stress that evaluations like ours and those of others do not serve as metrics for LLMs to beat before being deployed in high-stakes situations. We discourage gaming these benchmarks, and instead intend for this study to, like psychology studies, shed light on LLM behavior.
Acknowledgements
We would like to thank Tejas Vaidhya for contributing to experiments during the initial phase of the project. This work is supported by CIFAR and enabled in part by resources from Digital Research Alliance of Canada (https://alliancecan.ca), Mila (https://mila.quebec), and NVIDIA.
Appendix A Experiment Setup Details
A.1 Dataset Generation
Random ($\mathcal{D}_{1}$).
For causal graphs in this benchmark, the graph nodes are named as some arbitrary string of characters (E.g. ’xyz’) that are independently generated for each graph.
Tübingen ($\mathcal{D}_{2}$).
For the Tübingen benchmark, we select the graph nodes from the Tübingen pairs (TP) dataset in the following manner:
  • Bivariate. We randomly sample cause-effect pairs from the TP dataset and each cause-effect pair is used to generate a graph.
    Example - Altitude (A) - Temperature (B) is a cause-effect pair in the TP dataset, and the input graph states: Altitude causes temperature.
  • Confounding. We randomly sample two different cause effect pairs such that both have the common cause.
    Example - Altitude (A) - Temperature (B) and Altitude (A) - Precipitation (C) are cause-effect pairs in the TP dataset, and the input graph states: Altitude causes temperature ; Altitude causes precipitation.
  • Mediation. We randomly sample two different cause effect pairs such that the effect of one pair is the cause in the other pair.
    Example - Altitude (A) - Temperature (B) and Temperature (B) - Ozone concentration (C) are cause-efffect pairs in the TP dataset, and the input graph states: Altitude causes temperature ; Temperature causes Ozone concentration.
Anti-commonsense ($\mathcal{D}_{3}$).
Similar to the Tübingen ($\mathcal{D}_{2}$) case, we select the graph nodes from the TP dataset. However, we change the input causal relations between them in an adversarial manner as follows:
  • Bivariate. We randomly sample a cause-effect pairs from the TP dataset and the provide the reversed cause-effect relationship in the input graph.
    Example - Altitude (A) - Temperature (B) is a cause-effect pair in the TP dataset, and the input graph states: Temperature causes altitude.
  • Confounding. For variable $B$ and $C$, we first randomly sample cause effect pairs and then for sampled pair we randomly select a variable (for $A$) from the TP dataset that it is different from the variables in the corresponding cause-effect pair.
    Example - Altitude (B) - Temperature (C) is the sampled cause-effect pair and Age (A) is the randomly selected variable from the TP dataset. The relationships in the input graph are as follows: Age causes altitude ; Age causes temperature.
  • Mediation. For variable $A$ and $C$, we first randomly sample cause effect pairs and then for sampled pair we randomly select a variable (for $B$) from the TP dataset that it is different from the variables in the corresponding cause-effect pair.
    Example - Altitude (A) - Temperature (C) is the sampled cause-effect pair and Age (B) is the randomly selected variable from the TP dataset. The relationships in the input graph are as follows: Altitude causes age ; Age causes temperature.
A.2 Details regarding LLMs
Querying LLMs.
To query the GPT models, we used the OpenAI API  [^5] https://openai.com/blog/openai-api through the Langchain interface  [^6] https://python.langchain.com/docs/expression_language/interface. Meanwhile, Hugging Face’s Transformers API [^7] https://huggingface.co/docs/transformers/en/index was used to load and query LLaMA-2.
Regarding compute resources, since the GPT models are hosted remotely by OpenAI, we were able to make API calls locally with minimal CPU usage. Conversely, for LLaMA-2, we first had to load the model onto a cluster equipped with a GPU (RTX 8000/32 GB RAM) before submitting our queries.
LLM Output.
To ensure that LLMs respond with a yes or no to queries, we formalized a response format requiring LLMs to encapsulate their yes or no answers within an <answer></answer> tag. If the LLM does not adhere to this format, we initiate a retry, instructing it to comply with the required format. After the first retry, we relax the response format requirements, expecting only a yes or no answer. If after 10 retries the LLM fails to meet this criterion, we mark the attempt as a failure and attribute zero accuracy on the intervention effect prediction task.
Model Ids.
The specific model-ids of the LLMs are:
  • GPT-3.5: gpt-3-turbo-16k
  • GPT-4: gpt-4
  • GPT-4-turbo: gpt-4-turbo
  • LLaMA-2: llama-2-7b
A.3 Substitution Task
Figure 3: Prompt design for intervention reasoning in the substitution task.
For this task, we first describe the concept of interventions to LLM and refer to them as some arbitrary string of characters (E.g. ’xyz’). We then follow the same prompt template (Figure 2) to test the intervention effect performance, except we never use the word intervention again and replace it with the chosen substitution word.
A.4 Ground Truth Intervention Effects
We provide the ground truth value for every intervention effect task in our analysis in Table 5, i.e., we provide the true change in the causal relations after interventions.
This should help the reader understand clearly what queries we considered for the different research questions in our analysis. For the case of RQ3, we partition the queries into two sets based on the value of ground truth IE being either 0 or 1. While for the case of RQ2, we considered both the queries for the Bivariate case (intervention on B); while for the Mediator case (intervention on either B or C) we only considered the query ($A\rightarrow C$).
Graph Intervention Questions Ground Truth IE
Bivariate A $A\to B$ 0
$B\to A$ 0
B $A\to B$ 1
$B\to A$ 0
Confounding A $A\to B$ 0
$A\to C$ 0
$B\to C$ 0
B $A\to B$ 1
$A\to C$ 0
$B\to C$ 0
C $A\to B$ 0
$A\to C$ 1
$B\to C$ 0
Mediation A $A\to B$ 0
$A\to C$ 0
$B\to C$ 0
B $A\to B$ 1
$A\to C$ 1
$B\to C$ 0
C $A\to B$ 0
$A\to C$ 1
$B\to C$ 1
Table 5: Ground truth intervention effects for all the scenarios (causal graph, intervention variable, causal relation) considered in our benchmark.
Appendix B Additional Results
B.1 Reproducing results on the Tübingen dataset
We reproduced the results of Kıcıman et al. (2023) on the Tübingen dataset as shown in Table 6. Since their task did not involve the LLM to reason about the effect of interventions, we term this task of inferring relations from the input graph in the prompt as relation retrieval.
Model Accuracy
GPT-3 (text-davinci-003) 0.80
GPT-3.5 0.89
GPT-4 0.96
Table 6: Relation Retrieval accuracy of GPT models on the Tübingen dataset.
B.2 RQ1: How accurate are LLMs at predicting the effects of interventions?
To understand how effective LLMs are at IE prediction, we focused on the Random ($\mathcal{D}_{1}$) benchmark in the main paper (Table 1), as we want to remove any distracting effect of semantically meaningful entities as graph nodes. We now present the same results on the Tübingen $(\mathcal{D}_{2})$ and the Anti-commonsense $(\mathcal{D}_{3})$ benchmark as well in Table 7.
We find that the results in both the cases are very similar to the case with the Random ($\mathcal{D}_{1}$) benchmark; GPT-4-turbo performs the best and GPT models outperform LLaMA-2 by a big margin.
Graph Type Bivariate Confounding Mediation
Intervened Variable A B A B C A B C
GPT-3.5 $\mathcal{D}_{1}$ $0.83\pm 0.08$ $0.87\pm 0.06$ $0.8\pm 0.09$ $0.69\pm 0.12$ $0.36\pm 0.09$ $0.58\pm 0.11$ $0.36\pm 0.12$ $0.67\pm 0.12$
$\mathcal{D}_{2}$ $0.87\pm 0.06$ $0.83\pm 0.06$ $0.82\pm 0.08$ $0.76\pm 0.07$ $0.64\pm 0.07$ $0.67\pm 0.12$ $0.36\pm 0.09$ $0.67\pm 0.12$
$\mathcal{D}_{3}$ $0.57\pm 0.1$ $0.63\pm 0.09$ $0.82\pm 0.05$ $0.67\pm 0.11$ $0.31\pm 0.07$ $0.64\pm 0.08$ $0.42\pm 0.1$ $0.67\pm 0.12$
GPT-4 $\mathcal{D}_{1}$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.78\pm 0.09$ $0.82\pm 0.08$ $0.96\pm 0.03$
$\mathcal{D}_{2}$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
$\mathcal{D}_{3}$ $0.9\pm 0.07$ $0.97\pm 0.03$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $0.84\pm 0.07$ $1.0\pm 0.0$
GPT-4-turbo $\mathcal{D}_{1}$ $1.0\pm 0.0$ $0.97\pm 0.03$ $0.96\pm 0.03$ $0.93\pm 0.05$ $1.0\pm 0.0$ $0.98\pm 0.02$ $1.0\pm 0.0$ $1.0\pm 0.0$
$\mathcal{D}_{2}$ $0.93\pm 0.04$ $0.9\pm 0.05$ $0.87\pm 0.04$ $0.8\pm 0.08$ $1.0\pm 0.0$ $0.96\pm 0.03$ $1.0\pm 0.0$ $1.0\pm 0.0$
$\mathcal{D}_{3}$ $0.6\pm 0.11$ $0.57\pm 0.11$ $0.62\pm 0.09$ $0.78\pm 0.08$ $0.96\pm 0.03$ $0.69\pm 0.09$ $0.96\pm 0.04$ $0.96\pm 0.04$
LLaMA-2 $\mathcal{D}_{1}$ $0.5\pm 0.12$ $0.4\pm 0.12$ $0.56\pm 0.11$ $0.53\pm 0.11$ $0.16\pm 0.06$ $0.69\pm 0.09$ $0.56\pm 0.12$ $0.64\pm 0.12$
$\mathcal{D}_{2}$ $0.4\pm 0.12$ $0.2\pm 0.08$ $0.47\pm 0.12$ $0.38\pm 0.12$ $0.2\pm 0.07$ $0.58\pm 0.11$ $0.33\pm 0.12$ $0.67\pm 0.12$
$\mathcal{D}_{3}$ $0.33\pm 0.12$ $0.2\pm 0.06$ $0.42\pm 0.11$ $0.4\pm 0.11$ $0.09\pm 0.04$ $0.47\pm 0.12$ $0.31\pm 0.12$ $0.44\pm 0.1$
Table 7: IE prediction accuracy on all three $(\mathcal{D}_{1},\mathcal{D}_{2}\ \&\ \mathcal{D}_{2})$ benchmarks. GPT-4 variants are the best performing models, while LLaMA-2 appears to struggle with interventional reasoning.
B.3 RQ2: To what extent is LLM performance affected by memorized causal relations?
We provide results for the analysis to test for memorization with LLaMA-2 as well in Table 8. The performance of LLaMA-2 drops for the Mediation graph under intervention on variable C on the Anti-commonsense benchmark as compared to the Tübingen benchmark; indicating the negative impact of memorization on IE prediction.
We also add results on the Random benchmark, as another comparison point with results on the Anti-commonsense benchmark. The finding are similar to the case of comparing performance on the Anti-commonsense benchmark against the Tübingen benchmark (Table 2); the performance drops down even for the best performing models like GPT-4 & GPT-4-turbo.
Model Graph Intervened Variable Random ($\bf\mathcal{D}_{1}$) Tübingen ($\bf\mathcal{D}_{2}$) Anti-commonsense ($\bf\mathcal{D}_{3}$)
GPT-3.5 Bivariate B $0.87\pm 0.09$ $0.83\pm 0.1$ $0.63\pm 0.12$
Mediation B $0.4\pm 0.13$ $0.33\pm 0.12$ $0.4\pm 0.13$
C $0.67\pm 0.12$ $0.67\pm 0.12$ $0.67\pm 0.12$
GPT-4 Bivariate B $1.0\pm 0.0$ $1.0\pm 0.0$ $0.97\pm 0.05$
Mediation B $0.8\pm 0.1$ $1.0\pm 0.0$ $0.73\pm 0.11$
C $0.93\pm 0.06$ $1.0\pm 0.0$ $1.0\pm 0.0$
GPT-4-turbo Bivariate B $0.97\pm 0.05$ $0.9\pm 0.08$ $0.57\pm 0.13$
Mediation B $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
C $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
LLaMA-2 Bivariate B $0.4\pm 0.13$ $0.2\pm 0.1$ $0.2\pm 0.1$
Mediation B $0.47\pm 0.13$ $0.33\pm 0.12$ $0.33\pm 0.12$
C $0.6\pm 0.13$ $0.67\pm 0.12$ $0.4\pm 0.13$
Table 8: IE prediction performance on specific scenarios to understand the role memorization. Even best performing models in Table 1 like GPT-4 are negatively impacted by memorization effects in the Mediation scenario.
B.4 RQ3: Could LLMs be learning a shortcut to predict intervention effects?
For this analysis, as stated in the main paper, we only consider models that achieve an accuracy $\geq 0.95$ on relation retrieval. Hence, we first report in Table 9 the relation retrieval performance, the ability of LLM to infer relations from the input graph in the prompt. Since the LLaMA-2 model does not satisfy our constraint, we drop it from further analysis.
Model Bivariate Confounding Mediation
GPT-3.5 $1.0\pm 0.0$ $1.0\pm 0.0$ $0.99\pm 0.01$
GPT-4 $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
GPT-4-turbo $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
LLaMA-2 $0.87\pm 0.06$ $0.63\pm 0.03$ $0.86\pm 0.01$
Table 9: Relation Retrieval accuracy of models on Random ($\mathcal{D}_{1}$) benchmark. The GPT models show good performance but LLaMA-2 performs poorly. Hence, as per our criteria we drop the LLaMA-2 model for analysis in RQ3.
B.5 RQ4: Are LLMs robust to descriptions of interventions in-context?
To understand how effective LLMs are at IE prediction under the substitution case, we focused on the Random ($\mathcal{D}_{1}$) benchmark in the main paper (Table 4), as we want to remove any distracting effect of semantically meaningful entities as graph nodes. We now present the same results on the Tübingen $(\mathcal{D}_{2})$ and the Anti-commonsense $(\mathcal{D}_{3})$ benchmark as well in Table 10.
We find that the results in both the cases are very similar to the case with the Random ($\mathcal{D}_{1}$) benchmark; GPT-4-turbo performs the best and GPT models outperform LLaMA-2 by a big margin. Further, contrasting these results against Table 7, we see that the performance generally drops down even for the best performing LLMs like GPT-4 and GPT-4-turbo.
Graph Type Bivariate Confounding Mediation
Intervened Variable A B A B C A B C
GPT-3.5 $\mathcal{D}_{1}$ $0.67\pm 0.12$ $0.83\pm 0.06$ $0.56\pm 0.11$ $0.58\pm 0.11$ $1.0\pm 0.0$ $0.62\pm 0.1$ $0.44\pm 0.1$ $0.4\pm 0.12$
$\mathcal{D}_{2}$ $0.67\pm 0.12$ $0.83\pm 0.06$ $0.56\pm 0.11$ $0.78\pm 0.07$ $0.89\pm 0.04$ $0.64\pm 0.08$ $0.49\pm 0.12$ $0.69\pm 0.12$
$\mathcal{D}_{3}$ $0.57\pm 0.11$ $0.63\pm 0.1$ $0.62\pm 0.09$ $0.82\pm 0.06$ $0.82\pm 0.04$ $0.56\pm 0.08$ $0.44\pm 0.11$ $0.73\pm 0.08$
GPT-4 $\mathcal{D}_{1}$ $0.97\pm 0.03$ $1.0\pm 0.0$ $0.89\pm 0.05$ $0.82\pm 0.07$ $1.0\pm 0.0$ $0.96\pm 0.03$ $0.76\pm 0.1$ $1.0\pm 0.0$
$\mathcal{D}_{2}$ $0.8\pm 0.09$ $0.9\pm 0.05$ $0.76\pm 0.1$ $0.82\pm 0.07$ $1.0\pm 0.0$ $0.93\pm 0.03$ $0.96\pm 0.03$ $1.0\pm 0.0$
$\mathcal{D}_{3}$ $0.83\pm 0.08$ $0.97\pm 0.03$ $0.67\pm 0.12$ $0.78\pm 0.09$ $1.0\pm 0.0$ $0.87\pm 0.05$ $0.84\pm 0.07$ $0.98\pm 0.02$
GPT-4-turbo $\mathcal{D}_{1}$ $0.9\pm 0.07$ $0.97\pm 0.03$ $0.96\pm 0.03$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
$\mathcal{D}_{2}$ $0.8\pm 0.08$ $0.93\pm 0.04$ $0.73\pm 0.1$ $0.78\pm 0.09$ $1.0\pm 0.0$ $0.8\pm 0.08$ $0.98\pm 0.02$ $1.0\pm 0.0$
$\mathcal{D}_{3}$ $0.57\pm 0.11$ $0.57\pm 0.11$ $0.69\pm 0.1$ $0.98\pm 0.02$ $1.0\pm 0.0$ $0.71\pm 0.1$ $0.89\pm 0.06$ $0.98\pm 0.02$
LLaMA-2 $\mathcal{D}_{1}$ $0.57\pm 0.1$ $0.37\pm 0.09$ $0.56\pm 0.1$ $0.49\pm 0.08$ $0.51\pm 0.07$ $0.62\pm 0.07$ $0.51\pm 0.07$ $0.38\pm 0.04$
$\mathcal{D}_{2}$ $0.3\pm 0.11$ $0.23\pm 0.08$ $0.56\pm 0.11$ $0.42\pm 0.1$ $0.31\pm 0.09$ $0.51\pm 0.11$ $0.22\pm 0.08$ $0.42\pm 0.08$
$\mathcal{D}_{3}$ $0.33\pm 0.12$ $0.2\pm 0.08$ $0.44\pm 0.11$ $0.36\pm 0.1$ $0.13\pm 0.05$ $0.36\pm 0.09$ $0.31\pm 0.11$ $0.38\pm 0.09$
Table 10: IE prediction accuracy on the all $(\mathcal{D}_{1},\mathcal{D}_{2}\ \&\ \mathcal{D}_{3})$ benchmarks for the substitution task. The performance of GPT-3.5 & GPT-4 is worse under substitution compared to the non-substitution case (Table 7), while GPT-4-turbo does not show significant change for the Random ($\mathcal{D}_{1}$) benchmark.